A nurse is preparing to administer amoxicillin and clavulanate potassium (Augmentin) 40 mg/kg/day PO divided in equal doses every 12 hr to a toddler who weighs 25 lb. Available is amoxicillin and clavulanate potassium suspension 200 mg/5 mL. How many mL should the nurse administer per dose? (Round the answer to the nearest tenth. Use a leading zero if applicable. Do not use a trailing zero.)
Answer #1

Dose = 40 mg/ kg/ day

weight =25/2.2

=12.33kg

dose = 40* 12.33

= 493mg/ day

Available dose 200 mg in 5 mldose = desired/ available * volume

=493/200*5

= 12.33

so 12.5 ml is the dose which should be given in two devideddises

6.5 ml twice a day